OpenStudy (msmr):

If x = a + bi and y = a - bi, solve x + 6y = 7 for x.

mathslover (mathslover):

\[\large{a+bi + 6(a-bi)=7}\] \[\large{a+bi+6(a)-6(bi)=7}\] \[\large{a+6a+bi-6bi=7}\]

mathslover (mathslover):

Can you try to add the like terms... ? @MSMR ?

OpenStudy (ash2326):

We have \[x=a+bi\] and \[y=a-bi\] \[x+6y=7 \] just plugin the value of x and y \[(a+bi)+6(a-bi)=7\] combine the like terms, what would you get @MSMR ?

OpenStudy (aravindg):

@MSMR please interact ! others are trying to help you get to the answer thanks

OpenStudy (msmr):

I was trying it out, sorry

mathslover (mathslover):

No problem MSMR take your time...

mathslover (mathslover):

But just reply so that we can believe that you r trying

OpenStudy (msmr):

so if you combine like terms and get 7a = 5bi = 7, I'm a lil confused on what to do next.

mathslover (mathslover):

\[\large{7a-5bi=7}\]

mathslover (mathslover):

did u mean this?

OpenStudy (msmr):

right

OpenStudy (msmr):

but then how do you solve for x, since it's not even in the equation anymore? do you put it in a + bi form and divide out (if that makes sense?)

mathslover (mathslover):

wait

OpenStudy (aravindg):

@MSMR let me knw if you need an alternate soln

mathslover (mathslover):

do you have options?

OpenStudy (msmr):

that might be helpful because i'm not sure what to do once I've combined like terms

OpenStudy (msmr):

no, that's the whole problem and I don't have the answers.

mathslover (mathslover):

ok... \[\large{x+6y=7}\] \[\large{x=7-6y}\] \[\large{x=7-6(a-bi)}\] \[\large{x=7-6a+6bi}\]

mathslover (mathslover):

wht u think @AravindG and @ash2326

OpenStudy (ash2326):

@MSMR a complex no. is of the form a+bi a= real part b= imaginary part we got \[7a-5bi=7\] or we have \[7a-5bi=7+0i\] now compare both sides equating real part \[7a=7\] \[-5b=0\] can you find a and b now?

OpenStudy (ash2326):

*equating complex parts \[-5b=0\]

OpenStudy (msmr):

so is a=7 and b=0?

OpenStudy (ash2326):

Yes, you're right

OpenStudy (msmr):

so would x = 7? (7 + 0)

OpenStudy (ash2326):

yeah :)

OpenStudy (ash2326):

@MSMR do you get this?

OpenStudy (msmr):

I think so, yeah. I'll try to do the rest of my homework and see. Thank you (all of you!)

OpenStudy (ash2326):

good:D you're welcome :)